OpenStudy (anonymous):

The derivative of y=(x+x^-1)^2 and why I can't use chain rule?

hartnn (hartnn):

why do you think you can use chain rule ? you can.

OpenStudy (anonymous):

oh. well, when I tried to use it, I got a wrong answer, a really complicated one...

hartnn (hartnn):

then try expanding using \((a+b)^2=a^2+2ab+b^2\) this is a simpler way.

hartnn (hartnn):

a=x,b=x^-1

OpenStudy (anonymous):

really? hmmm I had my dad do the problem, and he did the chain rule, except without the du/dx, and he couldn't relly explain why... it seemed simpler, but I was confused on why he didn't put the du/dx~

hartnn (hartnn):

if chain rule is used, we need du/dx anyways, did you try expanding (x+1/x)^2 ?

OpenStudy (anonymous):

x^-1 times x^-1 is adding the exponents? verdad?

hartnn (hartnn):

\(x^{-1} \times x^{-1}=x^{-1-1}=..?\)

OpenStudy (anonymous):

okay, cool. I got 2(1-x^-3) when I expanded and took the derivative.

hartnn (hartnn):

umm..how ? before taking the derivative, did u get \(x^2+2+x^{-2}\) ??

OpenStudy (anonymous):

OH. whoops. 2x...? lol okay, I'll try it again.

hartnn (hartnn):

2(1-x^-3) is almost correct, just one term is in error

OpenStudy (anonymous):

when you factor out an x from something like (2x-2x^-3), what would you get for the second term in the paretheses?

hartnn (hartnn):

(2x-2x^-3), is correct answer, you can factor out 2 to get 2(x-x^-3) if you want to factor out x also, (which is not necessary, but since you've asked) \(2(x-x^{-3})=2(x-x \times x^{-4})=2x(1-x^{-4})\)

OpenStudy (anonymous):

okay, YES. wait how would you do this problem with the chain rule? I think I'll still need that for the test tomrorw...

hartnn (hartnn):

y= u^2 u =x+1/x dy/dx=...?

OpenStudy (anonymous):

2(x+1/x)=dy/dx?

hartnn (hartnn):

what ? you know the definition of chain rule, right ?

hartnn (hartnn):

\([f(g(x))]' = f'(g(x))g'(x)\)

hartnn (hartnn):

y=u^2 dy/dx = (u^2)' du/dx

hartnn (hartnn):

and du/dx=..?

hartnn (hartnn):

oh, you just did 2u and got 2(x+1/x) right ? you need to multiply it bu du/dx also

OpenStudy (anonymous):

so I need to multiply that by du/dx to get dy/dx???

hartnn (hartnn):

yes because, dy/dx = (dy/du) * (du/dx)

OpenStudy (anonymous):

So then idy/dx is basically like finding the derivative of the entire function and du/dx is like finding the derivative of only the inside?

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

So then that sepcific question that we were talking about earlier would become 2(x+1/x)(1+-1x^-2)?

hartnn (hartnn):

that would be correct :) good.

hartnn (hartnn):

if you simplify that, you would again gget (2x-2x^-3)

OpenStudy (anonymous):

I ended up at 2x+4/x+2/x^3...?

hartnn (hartnn):

you tried to simplify this: 2(x+1/x)(1-1x^-2) right ? how come there's no - *minus* sign :P

OpenStudy (anonymous):

OH!!! let me try again~ :)

hartnn (hartnn):

sure, take your time , but do it correctly this time :)

OpenStudy (anonymous):

2x-2/x^3?

hartnn (hartnn):

that.is.correct.:)

hartnn (hartnn):

which is same what we got...

OpenStudy (anonymous):

oh yeah~:) phfft I was totally thinking when I pu that post up...lol thank you!!!here, best response!!! <3 cya

hartnn (hartnn):

welcome ^_^ cya :)